OpenStudy (anonymous):

2x2 + 10x + 12 = 0 Solve using the Quadratic formula

OpenStudy (anonymous):

Do you know the Quadratic formula?

OpenStudy (anonymous):

x=-b+-sqrtb^2-4ac/2(a) ?

OpenStudy (***[isuru]***):

when \[ax^{2} + bx + c = 0\] \[x = \frac{ - b \pm \sqrt{ b^{2} - 4ac } }{ 2a }\] now in ur equation a = 2 b= 10 c= 12 plug the in to the equation can u do that ?

OpenStudy (anonymous):

x=-10sqrt1?

OpenStudy (***[isuru]***):

\[x = \frac{ -10 \pm \sqrt{ 10^{2} - 4(24) } }{ 4 }\]

OpenStudy (anonymous):

okay i got -10sqrt-1 what do i do now?

OpenStudy (***[isuru]***):

u will get \[x = \frac{ -10 \pm \sqrt{4} }{ 4 }\] which is \[x = \frac{ -10 + 2 }{ 4 } \ \ or \ x= \frac{ -10 -2 }{ 4 }\] so... \[x = -2 \ \ \ \ \ or \ \ \ \ \ x= -3\]

OpenStudy (***[isuru]***):

got it :) ?

OpenStudy (anonymous):

Yes THANKYOU!!!

OpenStudy (***[isuru]***):

u r welcome!